# Symmetric Preparation of Systems

### v2

@@ 2. Formal power series solutions
Having found $u^{*}$ we obtain $(u,m)$ by (4.3) in $\overline{\omega}$. This gives surjectivity and a tame right inverse to $d\mathcal{F}$ locally for any $(U,M)\in V$. By Theorem C.12, we obtain that the $C^{\infty}$ mapping $\mathcal{F}$ has a tame local right inverse on a neighborhood of $(F_{1},F_{0})$ in (4.1), giving a solution to (1.6). Observe that $\mathcal{F}$ is surjective also on some $(F_{1},F_{0})$ such that $F_{0}(0)\neq 0$ depending on $F_{1}$. Note that every antisymmetric $B$ gives a solution.
@@
It remains to prove uniqueness in the symmetric case for $(U,M)\in V$ when $U(t,x)>0$ and $u(t,x)$ is symmetric. We shall then show that there is only one symmetric solution to (1.6), by finding the corresponding $B(t,x)$ in (4.5). By adding the equation $\operatorname{Im}u^{*}\equiv\operatorname{Im}(C_{0}/2)\equiv 0$ with $C_{0}$ given by (4.12), we find
(4.15) $\left\{\begin{aligned} &2u_{0}^{*}-C_{01}^{0}Mu_{0}^{*}-C_{02}^{0}R-(C_{03}B)_{0}-(U^{-1}B)_{0}\equiv C_{00}^{0}\\ &-tC_{01}^{1}Mu_{0}^{*}+2tu_{1}^{*}-tC_{02}^{1}R-t(C_{03}B)_{1}-t(U^{-1}B)_{1}\equiv tC_{00}^{1}\\ &R_{01}Mu^{*}+R_{3}R+R_{03}B\equiv-R_{00}\\ &\operatorname{Im}(C_{01}Mu_{0}^{*})+\operatorname{Im}(C_{02}R)+\operatorname{Im}(C_{03}B)+\operatorname{Im}(U^{-1}B)\equiv-\operatorname{Im}(C_{00})\end{aligned}\right$.
By dividing the second equation by $t$, this system has following matrix for $(u_{0}^{*},u_{1}^{*},R,B)$
(4.16) $\begin{pmatrix}2\operatorname{Id}_{N}-C_{01}^{0}M&0&-C_{02}^{0}&-(C_{03}\,\boldsymbol{\cdot})_{0}-(U^{-1}\,\boldsymbol{\cdot})_{0}\\ -C_{01}^{1}M&2\operatorname{Id}_{N}&-C_{02}^{1}&-(C_{03}\,\boldsymbol{\cdot})_{1}-(U^{-1}\,\boldsymbol{\cdot})_{1}\\ R_{01}M&0&R_{3}&R_{03}\\ \operatorname{Im}(C_{01}M\,\boldsymbol{\cdot})&0&\operatorname{Im}(C_{02}\,\boldsymbol{\cdot})&\operatorname{Im}(C_{03}\,\boldsymbol{\cdot})+\operatorname{Im}(U^{-1}\,\boldsymbol{\cdot})\end{pmatrix}$
where $C\,\boldsymbol{\cdot}$ is the map $E\to CE$. Thus $\operatorname{Im}(U^{-1}\,\boldsymbol{\cdot})$ is the map $E\mapsto\operatorname{Im}(U^{-1}E)$, which is invertible by Lemma A.1 because $U^{-1}>0$. Since $R_{3}$ also is invertible and $C_{03}=R_{03}=0$ when $M\equiv 0$, we find that (4.16) is invertible when $M\equiv 0$, and therefore when $\|M(x)\|\ll 1$. Thus, if the neighborhood $V$ is small enough, then we obtain a unique smooth solution $(u_{0}^{*},u_{1}^{*},R,B)$ to (4.15) near $(0,0)$ giving symmetric $u\equiv u_{0}+tu_{1}\equiv C_{0}/2$ with $R_{0}\equiv 0$. By (4.3) this gives a unique symmetric solution $(u,m)\in C^{\infty}$ to (4.2). Observe that this gives a tame map by Example C.4, and the neighborhood $\overline{\omega}$ only depends on the bounds on $(U,M)\in V$. Thus we obtain in the positive definite case that $d\mathcal{F}$ is a tame bijection locally for any positive definite $U$ and symmetric $M$ such that $(U,M)\in V$, so by Theorem C.12 we obtain that the mapping $\mathcal{F}$ is a tame bijection locally in this case, giving a unique solution to (1.6). ∎
@@ Appendix A A symmetry lemma
First we consider the case of constant matrices. To simplify, we define the symmetric product $S(U,A)=\frac{1}{2}(UA+AU)$, which is symmetric when $A$ is symmetric and antisymmetric when $A$ is antisymmetric, thus $A\mapsto S(U,A)$ preserves the symmetry. Since any $N\times N$ matrix can be written uniquely as a sum of a symmetric and an antisymmetric matrix, we find that $A\to S(U,A)$ is a bijection for any $N\times N$ matrix $A$. We also have that linear ON mappings $O$ preserves the symmetry, since $OAO^{-1}=OAO^{*}$.
Thus, by using an ON base of eigenvectors to $U$, we find that $U=\{\mu_{j}\delta_{jk}\}_{jk}$ with $\mu_{j}>0,\ \forall\,j$, and $A=\{\alpha_{jk}\}_{jk}$. If $A$ is symmetric then $\alpha_{kj}=\overline{\alpha}_{jk}$ $\forall\,jk$, and if $A$ is antisymmetric then $\alpha_{kj}=-\overline{\alpha}_{jk}$ $\forall\,jk$. In both cases, we find that
@@
Appendix B Symmetric Malgrange division
In this section, we shall prove a symmetric version of the Malgrange preparation theorem for systems, by modifying the proofs in the Appendix of [2] which generalized [5, Sect. 7.5] to systems. In what follows, we shall let $|A|=\det A$ be the determinant of $N\times N$ matrices $A$, and let $\|A\|$ be the matrix norm.
We are going to divide matrix valued analytic functions with symmetric matrix valued (first order) polynomials. We shall let $P(t,B)=t\operatorname{Id}_{N}+B$, where $B$ is symmetric $N\times N$ matrix, $\omega$} be an open set in $\mathbf{C}$ such that $|P(t,B)|\neq 0$ when $t\in\partial{\omega}\in C^{1}$. If $G(t)$ is analytic $N\times N$ system in $\overline{\omega}$, then we have
(B.1) $G(t)=Q(t)P(t,B)+R{\qquad}t\in{\omega}$
where
(B.2) $Q(t)=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(s-t)^{-1}\,ds{\qquad}t\in{\omega}$
is analytic and
(B.3) $R$ $=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(P(s,B)-P(t,B))(s-t)^{-1}\,ds$
$=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}\,ds$
is constant. Observe that $Q(t)$ and $R$ depend analytically on $B$. The remainder RR\/ is uniquely determined by the condition that $RP(t,B)^{-1}$ is analytic when $t\not\in{\omega}$. In fact, it is easy to see that (B.3) gives
$RP(t,B)^{-1}=-(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(s-t)^{-1}\,ds{\qquad}t\not\in{\omega}$
which is analytic in $\complement{\omega}$. Then the uniqueness of RR\/ follows, since (B.1) implies that
$Q(t)=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(s-t)^{-1}\,ds-(2{\pi}i)^{-1}\int_{\partial{\omega}}RP(s,B)^{-1}(s-t)^{-1}\,ds$
when $t\in{\omega}$. If $RP(s,B)^{-1}$ is analytic in $\complement{\omega}$ we may push the integration contour to infinity in the second integral. Since the integrand is $O(|s|^{-2})$ in the first integral we obtain (B.2). Now we want to make this division in a small complex strip around $\mathbf{R}$, with uniform bounds.
Lemma B.1.
Let $G(t)$ be a bounded analytic $N\times N$ system in $\{\,t\in\mathbf{C}:|\operatorname{Im}t|<{\varepsilon}\,\}$. Then for any symmetric $N\times N$ matrix $B$ we can find analytic $N\times N$ system $Q^{{\varepsilon}}(t,B,G)$ and $R^{{\varepsilon}}(B,G)$ depending linearly on $G(t)$, such that
(B.4) $G(t)=Q^{{\varepsilon}}(t,B,G)P(t,B)+R^{{\varepsilon}}(B,G){\quad}\text{if}{\quad}\|B\|<1$
We also have the following estimates
(B.5) $\|\partial_{t}^{{\alpha}}\partial_{B}^{{\beta}}Q^{{\varepsilon}}(t,B,G)\|\leq C_{{\alpha}{\beta}}M{\varepsilon}^{-{\alpha}-|{\beta}|-2}$
$\|\partial_{B}^{{\beta}}R^{{\varepsilon}}(B,G)\|\leq C_{{\beta}}M{\varepsilon}^{-|{\beta}|-1}$
where $M=\sup\{\,\|G(s)\|:|\operatorname{Im}s|\leq{\varepsilon}\,\}$ and the constants are independent of ${0<\varepsilon}\leq 1$.
Observe that for the general case when $B$ is not symmetric, we find that the estimates (4.5) have the factor $N$ in front of $|\beta|$, see [2, Lemma A.1].
Proof.
This will be an adaptation of the proof of Lemma A.1 in [2]. Let $A$ be symmetric $N\times N$ matrix, and assume $\|A\|<1$. This gives that 1<t<1-1<t<1 when $|P(t,A)|=0$ since then tw=Awtw=-Aw for some $w\neq 0$. Thus we have that
(B.6) $|P(t,A)|\neq 0\qquad\text{when $\operatorname{Im}t\neq 0$}$
Let $\omega$} be the rectangle $\{\,s:|\operatorname{Im}s|<{\varepsilon},|\operatorname{Re}s|<2\,\}$, $0<\varepsilon\leq 1$, and let ${\omega}_{t}$ be the union of this rectangle and a congruent one centered at $\operatorname{Re}t$. Then (B.1) is satisfied when $|\operatorname{Im}t|<{\varepsilon}$, with
(B.7) $Q(t)=(2{\pi}i)^{-1}\int_{\partial{\omega}_{t}}G(s)P(s,A)^{-1}(s-t)^{-1}\,ds$
and
(B.8) $R=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,A)^{-1}\,ds$
Since $P(t,A)$ is symmetric, we find from (B.6)
(B.9) $\|P(s,A)^{-1}\|\leq{\varepsilon}^{-1}\quad\text{on}\quad\partial{\omega}_{t}\bigcup\partial{\omega}$
In fact, if $\alpha_{j}\in(-1,1)$ are the eigenvalues of $A$ then the eigenvalues of $P(t,A)$ are $t+\alpha_{j}$, and $\|P(t,A)^{-1}\|\leq\max_{j}|t+\alpha_{j}|^{-1}\leq\varepsilon^{-1}$ when $|\operatorname{Im}t|\geq\varepsilon$ or $|\operatorname{Re}t|\geq 2$. This gives the estimates
(B.10) $\left\{\begin{aligned} &\|Q(t)\|\leq CM{\varepsilon}^{-2}\quad\text{when $|\operatorname{Im}t|<{\varepsilon}$}\\ &\|R\|\leq CM{\varepsilon}^{-1}\end{aligned}\right$.
Now we shall perturb the coefficients of $P$. Since $\|P(s,A)^{-1}\|\leq{\varepsilon}^{-1}$ on $\partial{\omega}\cup\partial{\omega}_{t}$, we find $\|P(s,B)^{-1}\|\leq c{\varepsilon}^{-1}$ there, if $\|B-A\|\leq c_{0}{\varepsilon}$ for some $c_{0}\ll 1$. In fact, the eigenvalues of symmetric matrices depends Lipschitz continuously on the matrix norms by the minimax principle. Thus, by using (B.7)–(B.8) for P(t,B)P(t,B)\/ we obtain (B.4) with $Q^{{\varepsilon}}=Q$ and $R^{{\varepsilon}}=R$ when $\|B-A\|\leq c_{0}{\varepsilon}$. Clearly, QQ\/ and RR\/ are analytic functions of $B$. We easily obtain the estimates (B.5) when $|\operatorname{Im}t|<{\varepsilon}$ and $\|B-A\|\leq c_{0}{\varepsilon}$.
Next, take a suitable partition of unity $\{\,{\Psi}_{j}(X)\,\}$ for the symmetric matrices, which can be parametrized by $\mathbf{R}^{N}\times\mathbf{C}^{N(N-1)/2}\cong\mathbf{R}^{N^{2}}$, for example the one in [5, Theorem 1.4.6], and put ${\Phi}_{j}(X)={\Psi}_{j}(X/c{\varepsilon})$. If $c$ is small enough (independently of $\varepsilon$}), we obtain $\|X-Y\|<c_{0}{\varepsilon}$ if $X$, $Y\in\operatorname{\rm supp}{\Phi}_{j}$, $\forall\,j$. By doing the above construction with $A_{j}\in\operatorname{\rm supp}{\Phi}_{j}$, we get the functions $Q^{{\varepsilon}}_{j}$ and $R^{{\varepsilon}}_{j}$. By taking $Q^{{\varepsilon}}=\sum_{j}{\Phi}_{j}Q^{{\varepsilon}}_{j}$ and $R^{{\varepsilon}}=\sum_{j}{\Phi}_{j}R^{{\varepsilon}}_{j}$, we obtain the result. ∎
The proof of Proposition A.2 in [2] now gives the following result.
Proposition B.2.
Let $G(t)\in\mathcal{S}(\mathbf{R})$ with values in $N\times N$ matrices and let $B$ be a symmetric $N\times N$ matrix. Then we can find smooth $N\times N$ systems $Q(t,B,G)$ and $R(B,G)$ depending linearly on $G(t)$, such that (B.4) holds with QQ\/ and RR\/ instead of $Q^{{\varepsilon}}$ and $R^{{\varepsilon}}$. Also, we have the following estimates
(B.11) $\|\partial_{t}^{{\alpha}}\partial_{B}^{{\beta}}Q(t,B,G)\|\leq C_{{\alpha}{\beta}}\int(\|G\|+\|G^{(\nu)}\|)\,dt{\quad}\nu={\alpha}+|{\beta}|+4$
$\|\partial_{B}^{{\beta}}R(B,G)\|\leq C_{{\beta}}\int(\|G\|+\|G^{(\nu)}\|)\,dt{\quad}\nu=|{\beta}|+3$
Observe that in the general case, when $B$ is not symmetric, the number of derivatives of $G$ in (B.11) grows as $N|\beta|$ so the estimates are not tame in general, see Appendix C. For completeness, we shall prove Proposition B.2.
Proof.
Choose a function $\psi\in C_{0}^{\infty}$ such that $0\leq\psi\leq 1$, $\psi$ has support where $|t|\leq 2$ and $\psi=1$ on [1,1][-1,1]. Let
(B.12) $G_{0}(t)=(2\pi)^{-1}\int\widehat{G}(\tau)\psi(\tau)\,d\tau$
(B.13) $G_{j}(t)=(2\pi)^{-1}\int\widehat{G}(\tau)\big(\psi(2^{-j}\tau)-\psi(2^{1-j}\tau)\big)\,d\tau\qquad j\in\mathbf{Z}_{+}$
Clearly, $G_{j}$ is analytic $N\times N$ matrix and that $G=\sum_{j=0}^{\infty}G_{j}$. We also have
(B.14) $|\tau^{\nu}\widehat{G}(\tau)|\leq\int|G(t)|+|G^{(\nu)}(t)|\,dt$
which gives
(B.15) $|G_{j}(t)|\leq C2^{j(1-\nu)}\int|G(s)|+|G^{(\nu)}(s)|\,ds\qquad\text{if $|\operatorname{Im}t|<\varepsilon_{j}=2^{-j}$}$
Thus we find from Lemma B.1 that
(B.16) $G_{j}(t)=Q^{{\varepsilon}_{j}}(t,B,G)P(t,B)+R^{{\varepsilon}_{j}}(B,G){\quad}\text{if}{\quad}\|B\|<1$
where
(B.17) $\|\partial_{t}^{{\alpha}}\partial_{B}^{{\beta}}Q(t,B,G)\|\leq C_{{\alpha}{\beta}}2^{-j(\nu-{\alpha}-|{\beta}|-3)}\int(\|G\|+\|G^{(\nu)}\|)\,dt$
$\|\partial_{B}^{{\beta}}R(B,G)\|\leq C_{{\beta}}2^{-j(\nu-|{\beta}|-2)}\int(\|G\|+\|G^{(\nu)}\|)\,dt$
With $\nu$ chosen as in (B.11) we find by summing up the convergent series, that
(B.18) $Q(t,B,G)=\sum_{j=0}^{\infty}Q^{{\varepsilon}_{j}}(t,B,G)\text{ and }R(t,B,G)=\sum_{j=0}^{\infty}R^{{\varepsilon}_{j}}(t,B,G)$
are smooth functions such that (B.4) and (B.11) hold. ∎
Now, we can prove the following symmetric version of the Malgrange division theorem.
Theorem B.3.
Let $G(t,x)\in C^{\infty}_{0}(\overline{\omega})$ with values in $N\times N$ matrices, where $\omega$ is a neighborhood of the origin in $\mathbf{R}\times\mathbf{R}^{n}$, then for any $M(x)\in C^{\infty}(\overline{\omega})$ with values in symmetric $N\times N$ matrices we find
(B.19) $G(t,x)=Q(t,x)P(t,M(x))+R(x)\qquad(t,x)\in\omega$
where QQ\/ and RR\/ are smooth $N\times N$ systems depending linearly on $G$ satisfying
(B.20) $\|\partial_{t}^{{\alpha}}\partial_{x}^{{\beta}}Q(t,x)\|\leq C_{{\alpha}{\beta}}\|G\|_{j}{\quad}j={\alpha}+|{\beta}|+4$
$\|\partial_{x}^{{\beta}}R(x)\|\leq C_{{\beta}}\|G\|_{k}{\quad}k=|{\beta}|+3$
Here $\|F\|_{j}=\sum_{|\alpha|+|\beta|\leq j}\|\partial_{t}^{{\alpha}}\partial_{x}^{{\beta}}F\|$ is the $L^{2}$ Sobolev norm.
Observe that Theorem B.3 gives that the mapping
$(Q(t,x),R(x))\mapsto G(t,x)=Q(t,x)P(t,M(x))+R(x)$
is a $C^{\infty}$ surjective tame mapping, having a $C^{\infty}$ tame right inverse of degree 4, see Appendix C, but that mapping is not tame for Malgranges preparation theorem in general. Observe that the symmetric scalar case when $N=1$ follows directly from the inverse function theorem.
Proof.
If $G(t,x)\in C^{\infty}_{0}(\mathbf{R}\times\mathbf{R}^{n})$, then by using Proposition A.2 we obtain (B.19) with $Q(t,x)=Q(t,M(x),G(\cdot,x))$ and $R(x)=R(M(x),G(\cdot,x))$. Since $Q(t,x)$ and $R(x)$ depend linearly on $G(t,x)$ we obtain (B.20) from (B.11) by Leibniz’ rule, since any $L^{1}$ Sobelev norm of test functions is bounded by the corresponding $L^{2}$ norm. ∎
In the preparation theorems, we have only done right symmetric division by matrix valued polynomials. By taking transposes we also obtain the results for left symmetric divisions, this does not affect the norms of the matrices.
Corollary B.4.
Let $G(t,x)\in C^{\infty}_{0}(\overline{\omega})$ with values in $N\times N$ matrices, where $\omega$ is a neighborhood of the origin in $\mathbf{R}\times\mathbf{R}^{n}$, then for any $M(x)\in C^{\infty}(\overline{\omega})$ with values in symmetric $N\times N$ matrices we can write
(B.21) $G(t,x)=P(t,M(x))Q(t,x)+R(x)\qquad(t,x)\in\omega$
where QQ\/ and RR\/ are smooth $N\times N$ systems depending linearly on $G$ satisfying (B.20).
We find that (B.21) gives a $C^{\infty}$ surjective tame map of degree 0 having a $C^{\infty}$ tame right inverse of degree 4, see Example C.13.
Appendix C Tame Fréchet spaces and the Nash–Moser theorem

### v3

@@ 2. Formal power series solutions
Thus, we obtain a smooth solution ${\bf u}^{*}(t)\equiv T_{m}({\bf u}^{*})(t)+t^{m}{\bf u}^{*}_{1}(t)\equiv C_{0}(t)/2$ for $\mathbb{R}_{0}(t)\equiv 0$ in a sufficiently small compact neighborhood $\overline{\omega}$ of the origin. Observe that the neighborhood $\overline{\omega}$ only depends on the bounds on $U$, $\mathbb{M}$} and $B$. Having found ${\bf u}^{*}(t)$ satisfying (4.5) we obtain $({\bf u}(t),{\bf m}(t))$ by (4.4) in $\overline{\omega}$. This gives surjectivity and a tame local right inverse of degree $3+m$ to $d\mathcal{F}$ in $\overline{\omega}$ for any $(U(t),{\mathbb{M}}(t))\in V$. Observe that different antisymmetric $B$ gives different inverses.
@@
It remains to prove uniqueness in the symmetric case for $(U,{\mathbb{M}})\in V$, when $U$ and ${\bf u}(t)$ are symmetric and $U>0$ which then holds in $V$. We shall show that there is a unique symmetric solution to the equation, and find the corresponding $B(t)$ in (4.6). By adding the equation $\operatorname{Im}{\bf u}(t)\equiv\operatorname{Im}(C_{0})(t)\equiv 0$ with $C_{0}(t)$ given by (4.15) we find as before
(4.19) $\left\{\begin{aligned} &2T_{m}({\bf u})-T_{m}(C_{01}T_{m}({\mathbb{M}}{\bf u}^{*}))-T_{m}(C_{3}\mathbb{R})-T_{m}(C_{4}B)+T_{m}(U^{-1}B)\equiv T_{m}(C_{00})\\ &-C_{01}^{1}T_{m}({\mathbb{M}}{\bf u}^{*})+2{\bf u}_{1}-(C_{3}\mathbb{R})_{1}-(C_{4}B)_{1}-(U^{-1}B)_{1}\equiv C_{00}^{1}\\ &\mathbb{R}_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+\mathbb{R}_{3}\mathbb{R}+\mathbb{R}_{4}B\equiv-\mathbb{R}_{00}\\ &\operatorname{Im}(C_{01}T_{m}({\mathbb{M}}{\bf u}^{*}))+\operatorname{Im}(C_{3}\mathbb{R})+\operatorname{Im}(C_{4}B)+\operatorname{Im}(U^{-1}B)\equiv-\operatorname{Im}(C_{00})\end{aligned}\right$.
which is invertible near $t=0$ when $\|{\mathbb{M}}\|\ll 1$. In fact, the matrix of the system is
(4.20) $\begin{pmatrix}2\operatorname{Id}_{N}&0&-T_{m}(C_{3}\ \boldsymbol{\cdot})&-T_{m}(U^{-1}\ \boldsymbol{\cdot})\\ 0&2\operatorname{Id}_{N}&-(C_{3}\ \boldsymbol{\cdot})_{1}&-(U^{-1}\ \boldsymbol{\cdot})_{1}\\ 0&0&\mathbb{R}_{3}&0\\ 0&0&\operatorname{Im}(C_{3}\ \boldsymbol{\cdot})&\operatorname{Im}(U^{-1}\ \boldsymbol{\cdot})\end{pmatrix}$
modulo terms having local $C^{k}$ norms that are $\mathcal{O}(\|{\mathbb{M}}\|)$. Here $\operatorname{Im}(U^{-1}\ \boldsymbol{\cdot})$ is invertible on antisymmetric matrices by Lemma A.1. Since $\mathbb{R}_{3}(0)$ also is invertible, we find that (4.20) is invertible near the origin. As before, we find that (4.19) is invertible in a sufficiently small compact neighborhood $\overline{\omega}$ of the origin if $\|{\mathbb{M}}\|\ll 1$, which can be obtained by shrinking $V$ and $\overline{\omega}$. Thus, we obtain a unique smooth solution $({\bf u}(t),\mathbb{R}(t),B(t))$ to (4.19) with symmetric $\bf u$}, and by (4.4) a unique smooth symmetric solution $({\bf u}(t),{\bf m}(t))$ to (4.2).
Thus, $d\mathcal{F}$ is a tame bijection in a compact neighborhood $\overline{\omega}$ of the origin for any symmetric $(U(t),{\mathbb{M}}(t))\in V$. Observe that this gives a tame inverse $d\mathcal{F}^{-1}$ of degree $3+m$ by Example B.14, and the neighborhood $\overline{\omega}$ only depends on the bounds on $U$ and $\mathbb{M}$}. ∎
@@ Appendix A A symmetry lemma
First we consider the case of constant matrices. To simplify, we define the symmetric product $S(U,A)=\frac{1}{2}(UA+AU)$, which is symmetric when $A$ is symmetric and antisymmetric when $A$ is antisymmetric, thus $A\mapsto S(U,A)$ preserves the symmetry. We also have that linear ON mappings $O$ preserves the symmetry, since $OAO^{-1}=OAO^{*}$.
Next, we are going to prove that $A\mapsto S(U,A)$ is injective for both symmmetric and antisymmetric $A$. Since any $N\times N$ matrix can be written uniquely as a sum of a symmetric and an antisymmetric matrix, we will then find that $A\to S(U,A)$ is a bijection for any $N\times N$ matrix $A$. Thus, by using an ON base of eigenvectors to $U$, we find that $U=\{\mu_{j}\delta_{jk}\}_{jk}$ with $\mu_{j}>0,\ \forall\,j$, and $A=\{\alpha_{jk}\}_{jk}$. If $A$ is symmetric then $\alpha_{kj}=\alpha_{jk}$ $\forall\,jk$, and if $A$ is antisymmetric then $\alpha_{kj}=-\alpha_{jk}$ $\forall\,jk$. In both cases, we find that
@@
Appendix B Tame Fréchet spaces and the Nash–Moser theorem